# Bernoulli shifts on additive categories and algebraic $K$-theory of wreath products

### v2

@@ 4. $K$-theory of wreath products
To see that $\mathcal{A}\rtimes G$ is Noetherian, let $N\subseteq M$ be a submodule of a finitely generated $\mathbb{Z}(\mathcal{A}\rtimes G)$-module. Then $\iota^{*}N\subseteq\iota^{*}M$ is finitely generated by Claim 2 and by the flatness of $\iota^{*}$. Since $\mathcal{A}$ is Noetherian by assumption, we find an epimorphism $F\twoheadrightarrow\iota^{*}N$ where $F$ is a finitely generated free $\mathbb{Z}\mathcal{A}$-module. Since $\iota_{*}$ is right exact and preserves finitely generated free modules (see the comments at the end of [BL22, Section 5.A]), we get an epimorphism $\iota_{*}F\twoheadrightarrow\iota_{*}\iota^{*}N\twoheadrightarrow N$, where $\iota_{*}F$ is a finitely generated $\mathbb{Z}(\mathcal{A}\rtimes G)$-module. This implies that $N$ is finitely generated and therefore that $\mathcal{A}\rtimes G$ is Noetherian.
To see that $M$ is also regular coherent, let $M\in\mathrm{Mod}\text{-}\mathbb{Z}(\mathcal{A}\rtimes G)$ be finitely generated. Since $\iota^{*}M\in\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{A}$ is finitely generated by Claim 2, we can find a finite projective resolution $P_{\bullet}\to\iota^{*}M$. Then $\iota_{*}P_{\bullet}\to\iota_{*}\iota^{*}M$ is a finite projective resolution of $\mathbb{Z}(\mathcal{A}\rtimes G)$-modules. Since $M$ is a split summand of $\iota_{*}\iota^{*}M$ by Claim 1, it has a finite projective resolution as well. ∎
@@
Let $H$ be a finite group such that $\frac{1}{|H|}\in R$ and let $\varphi\in\mathrm{Aut}(H)$ be an automorphism. Let $Z$ be a $G$-set, where $G=H\rtimes_{\varphi}\mathbb{Z}$. Let $F\colon\mathcal{A}\to\mathcal{B}$ be a pointed virtual equivalence of pointed $R$-additive categories. Assume further that $\mathcal{A}^{\otimes n}$ and $\mathcal{B}^{\otimes n}$ are regular for every $n\geq 1$ and that for every inclusion of finite $H$-invariant subsets $Y\subseteq Y^{\prime}\subseteq Z$, the functors
@@
We prove the lemma by reducing it to Theorem 3.1. Denote by $x\in G=H\rtimes_{\varphi}\mathbb{Z}$ the generator $\mathbb{Z}$ so that $xhx^{-1}=\varphi(h)$ for all $h\in H$. We denote the shift actions on $\mathcal{A}^{\otimes Z}$ and $\mathcal{B}^{\otimes Z}$ by $\alpha$ and $\beta$ respectively. We define an automorphism $\Phi\in\mathrm{Aut}(\mathcal{A}^{\otimes Z}\rtimes H)$ by $A\mapsto\alpha_{x}(A)$ on objects and by
@@
$\sum_{(h,n)\in H\rtimes_{\varphi}\mathbb{Z}}f_{h,n}(h,n)\mapsto\sum_{h\in\mathbb{Z}}\left(\sum_{h\in H}f_{h,n}h\right)n$
@@
is an isomorphism. By compatibility of $K$-theory with filtered colimits (see [Qui73]), we may even replace $Z$ by a finite $H$-set $Y$. But now the statement follows from Theorem 3.1. ∎
@@
the virtual equivalence from Example 2.11, viewed as a pointed virtual equivalence with respect to the rank one modules $R\in\mathcal{F}^{p}(R),R[H]\in\mathcal{F}^{p}(R[H])$. Then the induced map
@@
This category is regular by Lemma 4.1 and thus $\mathcal{F}^{p}(R[H])\otimes\mathcal{F}^{p}(R[H])$ is regular by [BL22, Lemma 6.4]. By induction, $\mathcal{F}^{p}(R[H])^{\otimes n}$ is regular for every $n\geq 1$. Note that direct sum decomposition $R[H]\cong R\oplus I_{R}[H]$ induces an equivalence [^5] This equivalence is different from the virtual equivalence $F$. of categories $\mathcal{F}^{p}(R[H])\simeq\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H])$. In particular, $(\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H]))^{\otimes n}\simeq\mathcal{F}^{p}(R[H])^{\otimes n}$ is regular for every $n\geq 1$ as well. Moreover, functors
@@
Equipping $\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H])$ with the rank one module $R\in\mathcal{F}^{p}(R)$ as a base point, we have an equivariant equivalence
@@ 5. Bernoulli shifts on semi-simple algebras
This section is devoted to the computation of $K_{*}((\mathcal{A}\otimes{M_{n}^{\otimes Z}})\rtimes G)$ where $G$ is a group, $\mathcal{A}$ is an $R$-linear category with $G$-action, $Z$ is a $G$-set and $M_{n}$ denotes the $R$-algebra of $n\times n$-matrices over $R$ for some $n\geq 1$. We use this to calculate the $K$-theory of $A^{\otimes Z}\rtimes G$ where $A=M_{k_{0}}\oplus\dotsb\oplus M_{k_{n}}$ is a semi-simple $R$-algebra that decomposes into a finite sum of matrix algebras. This is mainly done by analysing the $A(G)$-module structure of $K_{*}(\mathcal{A}\rtimes G)$ where $A(G)$ is the Burnside ring of $G$, given by the Grothendieck group of the monoid of isomorphisms classes of finite $G$-sets. The addition in $A(G)$ is given by disjoint union and the multiplication is given by cartesian product. The Swan ring $\mathrm{Sw}^{fr}(G;\mathbb{Z})$ of $G$ is the Grothendieck group of the monoid of isomorphism classes of $\mathbb{Z}[G]$-modules that are finitely generated and free as $\mathbb{Z}$-modules, with direct sum as addition and tensor product as multiplication. There is a canonical homomorphism $A(G)\to\mathrm{Sw}^{fr}(G;\mathbb{Z})$ given by $X\mapsto\mathbb{Z}[X]$.
@@
$\Phi\colon\operatorname{Idem}((\mathcal{A}\otimes\operatorname{End}_{R}(R[X]))\rtimes G)\simeq\operatorname{Idem}(\mathcal{A}\rtimes G).$
The induced map
@@
Denote by $\lambda\colon A(G)\to\operatorname{End}_{\mathbb{Q}}(A(G)\otimes\mathbb{Q})$ the representation given by left multiplication. Let $(H_{i})_{1\leq i\leq k}$ be a collection of subgroups of $G$ containing each conjugacy class exactly once. Then $([G/H_{i}])_{1\leq j\leq k}$ is a basis for $A(G)$ so that any element $x\in A(G)$ can uniquely be written as $x=\sum_{1\leq i\leq k}n_{i}[G/H_{i}]$ with $n_{i}\in\mathbb{Z}$. By reordering the $H_{i}$ according to the maximal lenght of any chain of proper subgroups contained in $H_{i}$, we may furthermore assume that $i\leq j$ whenever $H_{i}$ subconjugate to $H_{j}$. We make the following observations:
@@
Now the theorem follows by letting $\alpha$ and $\beta$ be suitable polynomials in $[n^{Z}]$. The equality $[n^{Z}]^{l}=n\cdot\alpha$ implies that all the coefficients of $\alpha$ are positive. In particular $\alpha$ can be represented by a $G$-set. ∎
@@
where the upper horizontal arrows and the vertical arrows are induced by $\operatorname{id}\otimes 1$, the lower horizontal arrows are induced by $(\operatorname{id}\otimes 1)\otimes(\operatorname{id}\otimes 1)$, and the diagonal arrows are the canonical isomorphisms. This diagram commutes up to equivalence of functors. Since $K$-theory preserves filtered colimits [Qui73] and is invariant under equivalence of functors, the colimit of the vertical functors induces an isomorphism on $K$-theory. This proves the lemma. ∎
@@
Now the lemma follows from Lemma 5.4. ∎
@@
(5.1) $K_{*}((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes G)\xrightarrow{\sim}K_{*}((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}}\otimes M_{n}^{\otimes Z})\rtimes G)$
@@
Since $K$-theory preserves filtered colimits [Qui73], we may assume that $Z$ is finite (the general case follows by taking the colimit over all finite subsets of $Z$). By Lemma 5.2, we can identify the map (5.1) with the map
@@
By Theorem 4.3 we may assume that $G$ is finite-by cyclic. If $G$ itself is finite, then the theorem follows from Lemmas 5.5 and 5.6 If $G$ is not finite, then $G$ is a semidirect product $H\rtimes_{\varphi}\mathbb{Z}$ for some finite group $H$. As in the proof of Lemma 4.2, we get canonical isomorphisms
@@
Note that $(\mathcal{A}\otimes M_{n}^{\otimes Z})\rtimes H$ is the increasing union of the subcategories $(\mathcal{A}\otimes M_{n}^{\otimes Y})\rtimes H$ where $Y$ ranges over all finite $H$-invariant subsets of $Z$. Each of these subcategories is regular by Lemma 4.1666Here we also use that regularity passes to matrix amplifications, which can be deduced for instance from [BL22, Lemma 6.4]. and the inclusions $(\mathcal{A}\otimes M_{n}^{\otimes Y})\rtimes H\to(\mathcal{A}\otimes M_{n}^{\otimes Y^{\prime}})\rtimes H$ for finite $H$-invariant subsets $Y\subseteq Y^{\prime}\subseteq Z$ are easily seen to be flat. Thus their union is regular coherent by [BL22, Lemma 11.2]. Combined with [BL22, Theorem 10.1], this even shows that $((\mathcal{A}\otimes M_{n}^{\otimes Z})\rtimes H)[\mathbb{Z}^{m}]$ regular coherent for any $m\geq 0$. The same argument show that $((\mathcal{A}\otimes M_{n}^{\otimes Z}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes H)[\mathbb{Z}^{m}]$ and $((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes H)[\mathbb{Z}^{m}]$ are regular coherent for any $m\geq 1$. In view of the decompositions (5.2), [BL22, Theorem 7.8] now reduces the problem to checking that the maps
@@
is pointed with respect to the rank one modules $M_{n}\oplus\dotsb\oplus M_{n}$ and $M_{n_{0}}\oplus\dotsb\oplus M_{n_{k}}$ if and only if the representing matrix $X$ satisfies
@@
By [CEKN22, Corollary 3.4] we can find such a matrix satisfying in addition $X\in\mathrm{SL}(k+1,\mathbb{Z})$. By writing the inverse of $X$ as a difference of two matrices with non-negative integer entries, we conclude that the functor (5.3) is a pointed virtual equivalence. Combining Lemmas 3.1 and 4.2 with Theorem 4.3, we conclude that the induced map
@@
is pointed with respect to the base points $0\oplus\dotsb\oplus 0\oplus R$ and $R\oplus\dotsb\oplus R$. Comparing the infinite tensor products with respect to these different base points and applying the combination of Theorem 3.1 and Lemma 4.2 with Theorem 4.3, we obtain an isomorphism

### v4

@@ 4. $K$-theory of wreath products
We now explain why $\iota_{*}$ is exact. Since $\iota$ is essentially surjective, it is enough to see its restriction $\iota^{\ast}\iota_{\ast}$ is exact. For any $\mathbb{Z}\mathcal{A}$-module $M$, we have a natural isomorphism
$(\iota^{\ast}\iota_{\ast}M)(A)\cong\oplus_{g\in G}M(\alpha_{g}(A))$
for any object $A$ in $\mathcal{A}$. This is obtained by sending
$x\otimes f_{g}g\in M(B)\otimes_{\mathbb{Z}}\mathrm{Hom}_{\mathcal{A}\rtimes G}(\iota(A),\iota(B))$
for $f_{g}\in\mathrm{Hom}_{\mathcal{A}}(\alpha_{g}(A),B)$ to
$f_{g}^{*}x\in M(\alpha_{g}(A))$
We leave it to the reader to verify that this is well-defined and natural, establishing the natural isomorphism $\iota^{\ast}\iota_{\ast}M\cong\oplus_{g\in G}M(\alpha_{g}-)$. With this in hand, it follows that $\iota_{\ast}$ is exact.
To see that $\mathcal{A}\rtimes G$ is Noetherian, let $N\subseteq M$ be a submodule of a finitely generated $\mathbb{Z}(\mathcal{A}\rtimes G)$-module. Then $\iota^{*}N\subseteq\iota^{*}M$ is finitely generated by Claim 2 and by the exactness of $\iota^{*}$. Since $\mathcal{A}$ is Noetherian by assumption, we find an epimorphism $F\twoheadrightarrow\iota^{*}N$ where $F$ is a finitely generated free $\mathbb{Z}\mathcal{A}$-module. Since $\iota_{*}$ is right exact (being a left adjoint) and preserves finitely generated free modules (being a left Kan extension, see also the comments at the end of [BL22, Section 5.A]), we get an epimorphism $\iota_{*}F\twoheadrightarrow\iota_{*}\iota^{*}N\twoheadrightarrow N$, where $\iota_{*}F$ is a finitely generated $\mathbb{Z}(\mathcal{A}\rtimes G)$-module. This implies that $N$ is finitely generated and therefore that $\mathcal{A}\rtimes G$ is Noetherian.
To see that $M$ is also regular coherent, let $M\in\mathrm{Mod}\text{-}\mathbb{Z}(\mathcal{A}\rtimes G)$ be finitely generated. Since $\iota^{*}M\in\mathrm{Mod}\text{-}\mathbb{Z}\mathcal{A}$ is finitely generated by Claim 2, we can find a finite projective resolution $P_{\bullet}\to\iota^{*}M$. Then $\iota_{*}P_{\bullet}\to\iota_{*}\iota^{*}M$ is a finite projective resolution of $\mathbb{Z}(\mathcal{A}\rtimes G)$-modules. Since $M$ is a summand of $\iota_{*}\iota^{*}M$ by Claim 1, it has a finite projective resolution as well. Indeed, such a projective resolution can be constructed by iteratively taking the pullback of the resolution $\iota_{*}P_{\bullet}\to\iota_{*}\iota^{*}M$, starting with the split inclusion $M\to\iota_{*}\iota^{*}M$. The exactness of the resulting sequence can be verified directly, pointwise.
@@
Let $H$ be a finite group such that $\frac{1}{|H|}\in R$ and let $\varphi\in\mathrm{Aut}(H)$ be an automorphism. Let $Z$ be a $G$-set, where $G=H\rtimes_{\varphi}\mathbb{Z}$. Let $F\colon\mathcal{A}\to\mathcal{B}$ be a based $R$-linear functor of based $R$-additive categories which is furthermore a virtual equivalence. Assume further that $\mathcal{A}^{\otimes n}$ and $\mathcal{B}^{\otimes n}$ are regular for every $n\geq 1$ and that for every inclusion of finite $H$-invariant subsets $Y\subseteq Y^{\prime}\subseteq Z$, the functors
@@
We prove the lemma by reducing it to Theorem 3.1. Denote by $x\in G=H\rtimes_{\varphi}\mathbb{Z}$ the generator of $\mathbb{Z}$ so that $xhx^{-1}=\varphi(h)$ for all $h\in H$. We denote the shift actions on $\mathcal{A}^{\otimes Z}$ and $\mathcal{B}^{\otimes Z}$ by $\alpha$ and $\beta$ respectively. We define an automorphism $\Phi\in\mathrm{Aut}(\mathcal{A}^{\otimes Z}\rtimes H)$ by $A\mapsto\alpha_{x}(A)$ on objects and by
@@
$\sum_{(h,n)\in H\rtimes_{\varphi}\mathbb{Z}}f_{h,n}(h,n)\mapsto\sum_{n\in\mathbb{Z}}\left(\sum_{h\in H}f_{h,n}h\right)n$
@@
is an isomorphism. By compatibility of $K$-theory with filtered colimits (see [LS14, Corollary 7.2]), we may even replace $Z$ by a finite $H$-set $Y$. But now the statement follows from Theorem 3.1. ∎
@@
the virtual equivalence from Example 2.11, viewed as a based virtual equivalence with respect to the rank one modules $R\in\mathcal{F}^{p}(R),R[H]\in\mathcal{F}^{p}(R[H])$. Then the induced map
@@
This category is regular by Lemma 4.1 and thus $\mathcal{F}^{p}(R[H])\otimes\mathcal{F}^{p}(R[H])$ is regular by [BL22, Lemma 6.4]. By induction, $\mathcal{F}^{p}(R[H])^{\otimes n}$ is regular for every $n\geq 1$. Note that the direct sum decomposition $R[H]\cong R\oplus I_{R}[H]$ induces an equivalence [^7] This equivalence is different from the virtual equivalence $F$. of categories $\mathcal{F}^{p}(R[H])\simeq\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H])$. In particular, $(\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H]))^{\otimes n}\simeq\mathcal{F}^{p}(R[H])^{\otimes n}$ is regular for every $n\geq 1$ as well. Moreover, the functors
@@
Equipping $\mathcal{F}^{p}(R)\oplus\mathcal{F}^{p}(I_{R}[H])$ with the rank one module $R\in\mathcal{F}^{p}(R)$ as a base point, we have an equivariant equivalence (c.f. the binomial formula)
@@ 5. Bernoulli shifts on semi-simple algebras
This section is devoted to the computation of $K_{*}((\mathcal{A}\otimes{M_{n}^{\otimes Z}})\rtimes G)$ where $G$ is a group, $\mathcal{A}$ is an $R$-linear category with $G$-action, $Z$ is a $G$-set and $M_{n}$ denotes the $R$-algebra of $n\times n$-matrices over $R$ for some $n\geq 1$. We use this to calculate the $K$-theory of $A^{\otimes Z}\rtimes G$ where $A=M_{k_{0}}\oplus\dotsb\oplus M_{k_{n}}$ is a semi-simple $R$-algebra that decomposes into a finite sum of matrix algebras. This is mainly done by analysing the $A(G)$-module structure of $K_{*}(\mathcal{A}\rtimes G)$ where $A(G)$ is the Burnside ring of $G$, given by the Grothendieck group of the monoid of isomorphism classes of finite $G$-sets. The addition in $A(G)$ is given by disjoint union and the multiplication is given by cartesian product. The Swan ring $\mathrm{Sw}^{fr}(G;\mathbb{Z})$ of $G$ is the $K_{0}$-group of the exact category of $\mathbb{Z}[G]$-modules that are finitely generated and free as $\mathbb{Z}$-modules, equipped with the tensor product as multiplication. Explicitely, $\mathrm{Sw}^{fr}(G;\mathbb{Z})$ is generated by finitely generated $\mathbb{Z}$-free $\mathbb{Z}[G]$-modules with the relation $[X]+[Z]=[Y]$ whenever $0\to X\to Y\to Z\to 0$ is a short exact sequence. There is a canonical homomorphism $A(G)\to\mathrm{Sw}^{fr}(G;\mathbb{Z})$ given by $X\mapsto\mathbb{Z}[X]$.
@@
Proof.
The main idea is to associate to a $\mathbb{Z}[G]$-module $M\cong\mathbb{Z}^{d}$ the functor $\mathcal{A}\rtimes G\to\mathcal{A}\rtimes G$ which maps an object $A$ to the object $\bigoplus_{i=1}^{d}A\eqcolon A\otimes_{\mathbb{Z}}M$ and a morphism $\sum_{g\in G}f_{g}g$ to the morphism $\sum_{g\in G}(f_{g}\otimes_{\mathbb{Z}}l_{g})g$ where $l_{g}\colon M\to M$ denotes the action by $g$. We refer to [BR07, Lemma 9.1] for the the details. ∎
@@
$\Phi\colon\operatorname{Idem}((\mathcal{A}\otimes\operatorname{End}_{R}(R[X]))\rtimes G)\simeq\operatorname{Idem}(\mathcal{A}\rtimes G)$
such that the induced map
@@
Denote by $\lambda\colon A(G)\to\operatorname{End}_{\mathbb{Q}}(A(G)\otimes\mathbb{Q})$ the representation given by left multiplication. Let $(H_{i})_{1\leq i\leq k}$ be a collection of subgroups of $G$ containing each conjugacy class exactly once. Then $([G/H_{i}])_{1\leq j\leq k}$ is a basis for $A(G)$ so that any element $x\in A(G)$ can uniquely be written as $x=\sum_{1\leq i\leq k}n_{i}[G/H_{i}]$ with $n_{i}\in\mathbb{Z}$. By reordering the $H_{i}$ according to the maximal length of any chain of proper subgroups contained in $H_{i}$, we may furthermore assume that $i\leq j$ whenever $H_{i}$ is subconjugate to $H_{j}$. We make the following observations:
@@
Now the proposition follows by letting $\alpha$ and $\beta$ be suitable polynomials in $[n^{Z}]$. The equality $[n^{Z}]^{l}=n\cdot\alpha$ implies that all the coefficients of $\alpha$ are positive. In particular $\alpha$ can be represented by a $G$-set. ∎
@@
where the upper horizontal arrows and the vertical arrows are induced by $\operatorname{id}\otimes 1$, the lower horizontal arrows are induced by $(\operatorname{id}\otimes 1)\otimes(\operatorname{id}\otimes 1)$, and the diagonal arrows are the canonical isomorphisms. This diagram commutes up to equivalence of functors. Since $K$-theory preserves filtered colimits (see [LS14, Corollary 7.2]) and is invariant under equivalence of functors, the colimit of the vertical functors induces an isomorphism on $K$-theory. This proves the lemma. ∎
@@
Observe moreover that there is a canonical isomorphism
$(M_{n}^{\otimes\mathbb{N}}\otimes\operatorname{End}_{R}(R[X])^{\otimes\mathbb{N}})\rtimes G\cong M_{n}^{\otimes\mathbb{N}}\otimes(\operatorname{End}_{R}(R[X])^{\otimes\mathbb{N}}\rtimes G).$
of $R$-linear categories. Now the lemma follows from Lemma 5.4. ∎
@@
(5.1) $K_{*}((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes G)\xrightarrow{\sim}K_{*}((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}}\otimes M_{n}^{\otimes Z})\rtimes G).$
@@
Since $K$-theory preserves filtered colimits (see [LS14, Corollary 7.2]), we may assume that $Z$ is finite (the general case follows by taking the colimit over all finite subsets of $Z$). By Lemma 5.2, we can identify the map (5.1) with the map
@@
By Theorem 4.3 we may assume that $G$ is finite-by cyclic. If $G$ itself is finite, then the theorem follows from Lemmas 5.5 and 5.6. If $G$ is not finite, then $G$ is a semidirect product $H\rtimes_{\varphi}\mathbb{Z}$ for some finite group $H$. As in the proof of Lemma 4.2, we get canonical isomorphisms
@@
Note that $(\mathcal{A}\otimes M_{n}^{\otimes Z})\rtimes H$ is the increasing union of the subcategories $(\mathcal{A}\otimes M_{n}^{\otimes Y})\rtimes H$ where $Y$ ranges over all finite $H$-invariant subsets of $Z$. Each of these subcategories is regular by Lemma 4.1888Here we also use that regularity passes to matrix amplifications, which can be deduced for instance from [BL22, Lemma 6.4]. and the inclusions $(\mathcal{A}\otimes M_{n}^{\otimes Y})\rtimes H\to(\mathcal{A}\otimes M_{n}^{\otimes Y^{\prime}})\rtimes H$ for finite $H$-invariant subsets $Y\subseteq Y^{\prime}\subseteq Z$ are easily seen to be flat. Thus their union is regular coherent by [BL22, Lemma 11.2]. Combined with [BL22, Theorem 10.1], this even shows that $((\mathcal{A}\otimes M_{n}^{\otimes Z})\rtimes H)[\mathbb{Z}^{m}]$ regular coherent for any $m\geq 0$. The same argument shows that $((\mathcal{A}\otimes M_{n}^{\otimes Z}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes H)[\mathbb{Z}^{m}]$ and $((\mathcal{A}\otimes M_{n}^{\otimes\mathbb{N}})\rtimes H)[\mathbb{Z}^{m}]$ are regular coherent for any $m\geq 1$. In view of the decompositions (5.2), [BL22, Theorem 7.8] now reduces the problem to checking that the maps
@@
is based with respect to the rank one modules $M_{n}\oplus\dotsb\oplus M_{n}$ and $M_{n_{0}}\oplus\dotsb\oplus M_{n_{k}}$ if and only if the representing matrix $X$ satisfies
@@
By [CEKN22, Corollary 3.4] we can find such a matrix satisfying in addition $X\in\mathrm{SL}(k+1,\mathbb{Z})$. By writing the inverse of $X$ as a difference of two matrices with non-negative integer entries, we conclude that the functor (5.3) is a based virtual equivalence. Combining Lemmas 3.1 and 4.2 with Theorem 4.3, we conclude that the induced map
@@
is based with respect to the base points $0\oplus\dotsb\oplus 0\oplus R$ and $R\oplus\dotsb\oplus R$. Comparing the infinite tensor products with respect to these different base points and applying the combination of Theorem 3.1 and Lemma 4.2 with Theorem 4.3, we obtain an isomorphism